40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Mar 10 16:21:11 2022, Radhika, DailyProgress, Cryo vacuum chamber, Insulated outer shield cooldown started 3/10 9x
    Reply  Mon Mar 14 14:04:54 2022, Radhika, DailyProgress, Cryo vacuum chamber, Insulated outer shield cooldown started 3/10 
Message ID: 2730     Entry time: Thu Mar 10 16:21:11 2022     Reply to this: 2732
Author: Radhika 
Type: DailyProgress 
Category: Cryo vacuum chamber 
Subject: Insulated outer shield cooldown started 3/10 

[Stephen, Radhika]

Our goal for this week's cooldown was to tape peek sheets fully around the outer shield lip, to leave no bare aluminum contact area with the cold plate. Secondly, we wanted to diagnose and mend the issue preventing the heater from outputting any power. The full procedure was:

1. We allowed the chamber to vent, unbolted the chamber lid and outer/inner shield lids.

2. We noticed that the solder joints between the heater body and its leads had debonded [Attachment 1].

    a. The suspension frame was taken out of the chamber and the test mass was removed from the frame.

    b. In doing so, we noticed that the varnish joining 1. the heater to cigarette paper and 2. cigarette paper to Si was debonding in certain areas, likely due to Aquadag not being fully removed from the test mass in the area of contact [Attachment 2].

3. We wrapped the copper leads a few times around the heater "wings" and re-applied solder [Attachments 3, 4].

4. We cleaned off aquadag from a greater area on the test mass and applied varnish to re-bond the heater [Attachments 5, 6].

    a. We let the varnish cure for ~2 days with a small weight on top.

5. The outer shield was removed from the chamber (without unbolting/removing the inner shield), and a single layer of peek sheet was taped the whole way around the bottom lip [Attachment 7].

6. We re-inserted the outer shield and passed the RTDs back through.

    a. We reattached a few RTD lead pins/sockets that had broken off in handling.

7. Lastly, we placed the test mass back into the suspension and into the chamber.

8. Close out [Attachments 8, 9]

The vacuum pump was engaged and the cryocooler was turned on at ~3:30PM.

Attachment 1: 82E812D8-946F-4F93-95C2-A62EC189A1E5_1_105_c.jpeg  58 kB  | Hide | Hide all
82E812D8-946F-4F93-95C2-A62EC189A1E5_1_105_c.jpeg
Attachment 2: CFDD543D-17FB-4F49-A62B-090A15E47821_1_105_c.jpeg  50 kB  | Hide | Hide all
CFDD543D-17FB-4F49-A62B-090A15E47821_1_105_c.jpeg
Attachment 3: B5D683AC-986A-49B9-A1C3-379851C877E5_1_105_c.jpeg  67 kB  | Hide | Hide all
B5D683AC-986A-49B9-A1C3-379851C877E5_1_105_c.jpeg
Attachment 4: E06C8191-8A7F-4917-A97A-80F9F38BA614.jpeg  186 kB  | Hide | Hide all
E06C8191-8A7F-4917-A97A-80F9F38BA614.jpeg
Attachment 5: 59F4BAD9-C23E-4881-AC82-AEADD05B3D45.jpeg  170 kB  | Hide | Hide all
59F4BAD9-C23E-4881-AC82-AEADD05B3D45.jpeg
Attachment 6: IMG_3160.jpeg  149 kB  | Hide | Hide all
IMG_3160.jpeg
Attachment 7: C8A0E8EB-EB18-4B76-BF0D-6A67A4831C77_1_105_c.jpeg  187 kB  | Hide | Hide all
C8A0E8EB-EB18-4B76-BF0D-6A67A4831C77_1_105_c.jpeg
Attachment 8: IMG_3167.jpeg  257 kB  | Hide | Hide all
IMG_3167.jpeg
Attachment 9: IMG_3168.jpeg  303 kB  | Hide | Hide all
IMG_3168.jpeg
ELOG V3.1.3-